OpenStudy (anonymous):

Terrence mows lawns during the summer. On the weekdays, he gets $41 per lawn. On the weekends, he gets $46 per lawn.If Terrance mowed 24 lawns and earned $1064, how many lawns did he mow on the weekdays and how many on the weekends?

OpenStudy (anonymous):

weekday lawn + weekend lawn = total $41 x + 46y = $1,064 @Educate_Me can you make the second equation to solve for x, y?

OpenStudy (anonymous):

x + y = 24 ?

OpenStudy (anonymous):

-> x = 24 - y => 41(24 - y) + 46y = $1,064 Solve for y, can you!